# Safe Trajectory Sets for Online Operation of Power Systems under Uncertainty

### v1

@@ Abstract
Flexibility provision from active distribution grids requires efficient and robust methods of optimization and control suitable to online operation. In this paper we introduce conditions for the safe operation of feedback optimization based controllers. We use the feasible operating region of a controlled system as bounds for safe system states and evaluate the trajectories of the controller based on the projection of the full system state onto the two-dimensional PQ-plane. We demonstrate the defined conditions for an exemplary sub-transmission system. We show that the proposed method is suitable to evaluate controller performance and robustness for systems subject to disturbances.
@@
Ancillary services, curative system operation, flexibility, online feedback optimization.
I Introduction
Distribution grids are increasingly taking on an active role in power system operations, including the provision of ancillary services by distributed energy resources (DER), aggregators, or distribution system operators (DSOs). This includes frequency control [1, 2], voltage support for the overlaying grid [3, 4, 5], and congestion management on transmission system level [7, 6]. The involvement of potentially large numbers of DERs and subsystems necessitates efficient and robust control and optimization methods tailored to the specific ancillary services. Given the high safety requirements in power systems, a resulting control architecture must be validated for performance and robustness, even under uncertainty, disturbances, and noise. To address these challenges of online control, there has been growing interest in the application of closed-loop techniques for solving optimization problems in power systems [8, 9]. These methods leverage the advantages of feedback control by incorporating real-time measurements while solving the optimization problem, leading to an iterative trajectory of systems states tracking the optimal solution. This approach offers several advantages, including robustness to model mismatch, resilience to disturbances, and lower computational complexity compared to conventional optimization problems such as the AC optimal power flow.
@@ I-A Related Work
Distribution grid flexibility is mainly discussed taking one of two perspectives. First, the quantification of available flexibility is of interest to allow for a consideration in operational planning [10]. Second the dispatch of a flexibility request requiring coordination of multiple, often numerous DER [11, 12].
@@
Power system optimization incorporating real-time feedback has been proposed for a variety of different use-cases. An approach to track the solution of the AC optimal power flow on the physical power system is presented in [20]. Approaches usually entail solving optimization problems by iteratively steering the physical system to an optimal solution based on linear input output sensitivities [21]. These can either be determined offline [22] or learned from measurement [23]. An approach to real-time coordination of DER on distribution grid level according to specified performance objectives, such as set point tracking at the substation is described in [24]. The authors of [25] describe an approach to voltage control on DSO level for different time scales. From a systemic perspective the application of Online Feedback Optimization (OFO) in provision of ancillary services requires the optimization and activation of such measures in online grid operation. An approach to curative grid operation using Online Feedback Optimization is introduced in [27]. OFO is performing supervisory control, recognizing critical system states and activating remedial actions on the same system layer if needed. To further leverage flexibility that is connected to underlying grid layers, i.e. sub-transmission and distribution systems, a hierarchical control structure for an efficient dispatch based on OFO for curative flexibility is introduced in [28, 29]. Different approaches to tuning the proposed control stack are introduced in [30] and [31]. These works illustrate, that the performance of a control stack based on OFO depends on the chosen formulation of the optimization problem and other parameters, e.g. the gain of the controller. The application in safety critical infrastructure like power systems therefore leads to a trade-off between robustness and performance of a proposed control stack. So far previous publications on flexibility quantification, flexibility coordination and feedback-based optimization lack methods to verify a specific control stack with respect to the problem of robustness vs. performance.
I-B Main Contribution
@@
In order address this gap the main contribution of this work is twofold: First, we introduce an extended hierarchical control architecture for multi-layer flexibility provision in online grid operation. To this end, we incorporate underlying grid layers as flexible resources to support load flow control. Second, we propose a novel method for calculating trajectory sets under uncertainty. This approach verifies the proposed control stack in terms of safety and robustness while quantifying the available flexibility based on the specific implementation and controller tuning. To achieve this, we systematically analyze trajectories resulting from adapting to new load flow set points at the point of common coupling between two grid layers. Based on the theoretical grantees and previous studies we expect OFO to ensure robust and safe operation of power systems under uncertainty and disturbances by leveraging real-time measurements to iteratively adapt control actions, even in scenarios with limited controllability and incomplete system models.
II Closed-Loop Control Architecture
@@
II-A Flexibility Provision
We denote $N$ as the set of busses, $B$ as the set of branches (i.e., lines and transformers), and $F$ as the set of all flexible actors in a grid. Let $\mathbf{v}\in\mathbb{R}^{|N|}$ represent the vector of bus voltages, $\mathbf{s}\in\mathbb{R}^{|B|}$ the vector of complex branch flows, and $\mathbf{p}\in\mathbb{R}^{|F|}$ and $\mathbf{q}\in\mathbb{R}^{|F|}$ the vectors of active and reactive power injections for flexible actors. For a given set point of apparent power flow at the PCC between two grid layers, $\mathbf{s}_{\text{set}}=[p_{\text{set}},q_{\text{set}}]^{T}$, we formulate the following dispatch problem for flexibility:
$\min_{\mathbf{v},\mathbf{s},\mathbf{p},\mathbf{q}}$ $\quad\Phi=||p_{\text{set}}-p_{\text{PCC}}||^{2}+||q_{\text{set}}-q_{\text{PCC}}||^{2}$ (1)
@@ II-A Flexibility Provision
This optimization minimizes the Euclidean norm of the difference between the measured apparent power flow, $\mathbf{s}_{\text{PCC}}=[p_{\text{PCC}},q_{\text{PCC}}]^{T}$, and the set point, $\mathbf{s}_{\text{set}}=[p_{\text{set}},q_{\text{set}}]^{T}$, to achieve the desired operating point while satisfying system technical constraints. These constraints, detailed in (1), include limits on bus voltages, $v_{n}\ \forall n\in N$, and branch flows, $s_{i}\ \forall i\in B$. Power limits for available flexibility, such as controllable loads, generation, and other grid layers, are also incorporated as inequality constraints. Since this optimization problem is solved online on the physical system, the AC power flow equations are not explicitly included as equality constraints. Instead, constraint satisfaction is ensured through real-time voltage and branch flow measurements. In the feedback optimization-based approach, the dispatch problem is solved iteratively. At each step, the cost function and constraints are incorporated into a quadratic optimization problem to determine the controller’s next action. The detailed formulation of the proposed OFO (Online Feedback Optimization) controller, including its iterative implementation, is presented in the following section.
II-B Feedback Optimization Based Control
In this section, we introduce the general concept of the OFO controller presented in this work. As stated in subsection I-A, dispatch problems in the form of (1) are typically solved offline for a model of the system with an estimate of possible disturbances. In contrast, the feedback-based approach iteratively steers the system by directly acquiring the relevant states through online measurements (see Figure 2). In this work we choose projected gradient descent as the underlying optimization algorithm for the controller.
@@ II-B Feedback Optimization Based Control
We consider a system $y\in\mathbb{R}^{p}$ described by its inputs $u\in\mathbb{R}^{q}$, a function mapping inputs to outputs $h(u)$ and disturbances $d\in\mathbb{R}^{p}$, where:
$y=h(u)+d$ (2)
For our presented use case, this entails a flexibility-providing power system that includes both controllable and non-controllable units, e.g. loads and DERs. The dispatch of flexibility is optimized based on (1) resulting in new operating points for flexible units. After actuation, all units adjust their feed-in or consumption of active and reactive power accordingly. To ensure stability and convergence to the optimum of the cost function and to allow for an analysis of steady-states, we separate the controller’s actions from the fast dynamics of the system [32]. This can be achieved by selecting an appropriate cycle time between the iterations awaiting all relevant faster dynamics of the system such as frequency and voltage transients. In this way we can abstract the differential-algebraic equations of the system describing its physical behavior to their resulting steady-states. We express this in the following assumption.
@@
To drive the system to the optimal solution of the underlying optimization problem (1) we approximate the behavior of the system by its steady-state map. The proposed controller utilizes an input-output sensitivity map $\nabla h$ that represents the gradient of system output $y$ with respect to control input $u$. The sensitivities are determined by linearly approximating around the smooth power flow manifold for an initial operating point. They include the relevant states, such as bus voltages, branch flows and especially power flow at the PCC. We calculate $\nabla h$ with:
@@
Due to the inclusion of feedback in each step of the controller, $\nabla h$ is commutable offline for a fixed initial operating point. We are subsequently able to use this constant approximation in online operation. For the controller to act on the system we define the vector of measurements $y\in\mathbb{R}^{p}$ with
@@
II-C Hierarchical Control Stack
In order to participate in ancillary services, the flexibility from DER at sub-transmission level needs to be aggregated at the TSO-DSO-interface, i.e. the coupling transformer between EHV and HV layer. In the approach presented in this paper, this leads to the following interactions between the supervisory control at TSO-level and the flexibility control at DSO-level. First, the flexibility potential must be estimated by the DSO and made known to the TSO. Second, the TSO needs to request flexibility by sending set points for power flow to the DSO, denoted as $\mathbf{s}_{set}$. To track the requested set points the controller on sub-transmission level needs to update its cost function $\Phi$ upon receiving a new set point (see (1) and (6)). This results in the following sequence of steps for each iteration of the control stack:
@@ II-C Hierarchical Control Stack
    Acquire the current grid $\mathbf{y}$ state by measurement of power flows and bus voltages.
@@
III Robust Trajectory Sets
@@
III-A Feasibility of System States
To evaluate the safety of systems states that result from actuation of a central controller we require a known safe set of such states considering the constraints of the system as a baseline to evaluate against.
@@ III-A Feasibility of System States
In this approach we maximize the load flow for a fixed direction in the solution space, subject to equality and inequality constraints of the AC-OPF problem. The FOR can subsequently be interpreted as a projection of the high-dimensional solution space of (10) onto the two-dimensional PQ-plane at the point of common coupling. Therefore, we are able to use the FOR $\mathcal{F}$ to validate individual system states $\mathbf{y}(k)$ that result from the actuation of the controller. As we choose the inequality constraints $g(\mathbf{x})$ in (1) to be identical to (10) we can make the following assumption for system states resulting from control actions during flexibility provision. We denote the projection of the full system state $\mathbf{y}$ onto the feasible operating region as $\mathbf{y}^{*}$.
@@
Using this assumption we are able to validate individual system states, system trajectories and consequently trajectory sets that result from the actions of the proposed controller. In the next section we define conditions for the safeness of trajectory sets with respect to the system constraints using the FOR as reference for safe operation.
III-B Feasibility of System Trajectories
@@
For the proposed formulations, safeness implies conditional safeness. Lastly we define unsafe trajectory sets based on constraint satisfaction of the final states of all trajectories $x(k)\in\mathcal{E}$.
@@
III-C Trajectory Sets under Uncertainty
@@
Using this definition we evaluate a controller stack for a fixed formulation of the optimization problem, a fixed gain $\alpha$ for a set of uncertainties and noise. We describe the transition of system states from $y_{k-1}$ to $y_{k}$ by the control actions $u(k-1)$ and by modeling the individual uncertainties where:
$y(k)=h^{*}(u)+\omega_{l}+\omega_{y}$ (15)
@@
In cases where we only assume controllability of generation units we consequently consider the current operating points of loads within the system as one of the main disturbances affecting the distribution grid during control of flexibility. We model the variation of the load over time according to [20] for a load covariance matrix $\Sigma_{l}$ with: $\omega_{l}\sim\mathcal{N}(0,\Sigma_{l})$. We differentiate the loads by type (households, industry and commercial) and adjust the entrances of the covariance matrix accordingly.
@@
As the proposed control architecture is using direct measurements of bus voltages and branch flows as feedback in the loop it is subject to any noise present in the respective signals. We model the noise term as uniform with $\omega_{y}\sim\mathcal{U}([a,b])$ uncorrelated for all measurements in $y$.
@@
Additionally to the non-modeled dynamics of the system, i.e. uncertainty of load and noisy measurements, we consider a mismatch in model and physical system by introducing bounded errors to the steady-state sensitivity map $\nabla h(u)$. These are described by a bounded uniform distribution $\omega_{s}\sim\mathcal{U}([c,d])$. This results in the augmented sensitivity map $\nabla h^{*}(u)$. Using the uncertainties and their respective distributions, we perform Monte Carlo simulations of the controller converging to reference values on the FOR. This allows for an evaluation of the performance and robustness of the controller in presence of disturbances. We are using the conditions we defined in subsection III-A to validate the resulting trajectory sets. Furthermore, we evaluate the iterations needed for convergence to the requested load flows. The results of these case studies are detailed in the next section.
IV Case Studies
@@ IV-A Scenario Description
IV-B Trajectory Sets at Point of Common Coupling
@@
IV-C Influence of Disturbance
This case study examines the behavior of trajectory sets under uncertainty and noisy measurements, focusing on the high-voltage sub-transmission grid described in subsection IV-A. The analysis evaluates trajectories that result from controlling load flow at the PCC to the transmission system. Parameter uncertainty, $\omega_{s}$, is modeled as a uniform distribution with bounds of ±$\pm 5\%$, while measurement noise, $\omega_{y}$, is similarly modeled with bounds of ±$\pm 2\%$ on all entries in $y$. First, we consider both load and generation of the system to be fully controllable. The controller is evaluated by performing in total 63.000 simulations of the trajectories with 500 iterations of OFO each under varying disturbances. Based on the resulting system states we can evaluate the density of the different load flows resulting from the control actions at the coupling transformers on $\mathcal{F}$ as shown in Figure 8.
@@ IV-C Influence of Disturbance
The density clearly reflects the general development of the trajectory sets for the controller as shown in Figure 7, indicating that even in when subject to uncertainties the control actions follow a clear trajectory. It is observable that the density is decreasing for later iterations of the controller. This is explainable by the increasing number of constraints that can become active as the system operates closer to its bounds. Figure 9 illustrates the density of system states lying outside the FOR.
@@
Approximately 40% of all trajectories within the resulting sets contain critical system states. By projecting the overall system states onto the FOR, we identify the limiting constraints under uncertainty. As shown in Figure 9, critical system states are primarily distributed along the lower right section of the FOR, beginning at $\underline{s}_{PCC}=-500\ \text{MW}-j600\ \text{MVAr}$. This distribution is attributed to the controller’s behavior in response to active limiting constraints during operation. Specifically, the constraints imposed by the PCC transformers, as described in subsection IV-A, limit the load flow in this region of the FOR. The observed distribution of critical states suggests that the controller struggles to accommodate the combined effects of parameter uncertainty and transformer constraints in the lower right section of the FOR. A detailed analysis of the sensitivity of the system states to $\omega_{s}$ and $\omega_{y}$ confirms that deviations in measurements significantly affect the load flow trajectory, especially near the constraint-dominated regions. As shown in Figure 10 the final trajectory sets under uncertainty vary significantly from the final set without disturbances.
@@
Especially, the achievable trajectory set of load flows for reactive power provision (top bound of $\mathcal{F}$) to the superimposed system is notably smaller than in subsection IV-B. This again can be attributed to the density of active constraints in that region of $\mathcal{F}$. As the individual upper bus voltage limits are nearly reached, the noisy measurement leads OFO to perceive these constraints as active. As the noise of the individual sensors is not correlated this leads to a non-convergence of trajectories as OFO is reacting simultaneously to varying perceived constraint violations within the range of the measurement noise $\omega_{y}$. This is further illustrated in Figure 11 showing an exemplary trajectory with and without noisy measurements.
@@
This highlights the difference in influence of measurement uncertainty and sensitivity mismatch as the controller without measurement noise is able to converge even with an augmented sensitivity map. This aligns with previous theoretical considerations guaranteeing convergence of projected gradient descent to an optimal solution for a sufficiently small gain $\alpha$ [26]. To further validate the proposed methods, we analyze a scenario without load controllability and present the corresponding final trajectory set in Figure 12. In this case, the controller attempts to converge to the vertices of the Feasible Operating Region (FOR) with a reduced resolution of $\mathcal{F}$.
@@
The reduced availability of controllable power is clearly reflected in the size of the trajectory set, which is significantly smaller compared to the scenario with full controllability of all actuators. The uncertain behavior of the loads introduces additional variability, as depicted in the final trajectory set (gray), diverging from the base scenario without uncertainty. Despite this, the control stack demonstrates robustness as defined in (14), with no observed violations of system limits across all evaluated states. This result highlights the adaptability of the proposed control framework under constrained controllability and uncertain system conditions. While the reduced flexibility imposes stricter limitations on achievable trajectory sets, the controller successfully maintains safe operation by dynamically adjusting to the available flexibility and uncertainties.
V Conclusion

### v2

@@ Abstract
Flexibility provision from active distribution grids requires efficient and robust methods of optimization and control suitable to online operation. In this paper we introduce conditions for the secure operation of power systems using feedback optimization based controllers. The feasibility of system operation is defined using the Feasible Operating Region (FOR), which serves as a boundary for safe system states. We propose a method to compute safe trajectory sets by projecting the high-dimensional system state onto the two-dimensional PQ-plane, enabling intuitive analysis of controller behavior. Our method is demonstrated on an exemplary sub-transmission system, where we evaluate the controller’s performance and robustness under disturbances and uncertainties. Results highlight the suitability of the proposed approach for assessing reachable and feasible system states in cases where the controlled grid is subjected to disturbances.
@@
Ancillary services, curative system operation, flexibility, online feedback optimization
Introduction
Distribution grids are increasingly taking on an active role in power system operations, including the provision of ancillary services by distributed energy resources (DER), aggregators, or distribution system operators (DSOs). This includes frequency control [1, 2], voltage support for the overlaying grid [3, 4, 5], and congestion management on transmission system level [6, 7]. The involvement of potentially large numbers of DERs and subsystems necessitates efficient and robust control and optimization methods tailored to the specific ancillary services. Given the high security requirements in power systems, a potential control architecture must be validated for performance and robustness, even under uncertainty, disturbances, and noise. To address these challenges of online control, there has been growing interest in the application of closed-loop techniques for solving optimization problems in power systems [8, 9]. These methods leverage the advantages of feedback control by incorporating real-time measurements while solving an optimization problem, leading to an iterative trajectory of systems states tracking the optimal solution. This approach offers several advantages, including robustness to model mismatch, resilience to disturbances, and lower computational complexity compared to conventional optimization problems such as the AC optimal power flow.
@@ Introduction
Distribution grid flexibility is mainly discussed taking one of two perspectives. First, the quantification of available flexibility is of interest to allow for a consideration in operational planning [10]. Second, the dispatch of a flexibility request requiring coordination of multiple, often numerous DER [11, 12].
@@
Power system optimization incorporating real-time feedback has been proposed for a variety of different use-cases. An approach to track the solution of the AC optimal power flow on the physical power system is presented in [20]. Approaches usually entail solving optimization problems by iteratively steering the physical system to an optimal solution based on linear input output sensitivities [21, 22]. These can either be determined offline [23] or learned from measurement [24]. An approach to real-time coordination of DER on distribution grid level according to specified performance objectives, such as set point tracking at the substation is described in [25]. The authors of [26] describe an approach to voltage control on DSO level for different time scales. From a systemic perspective the application of Online Feedback Optimization (OFO) in provision of ancillary services requires the optimization and activation of such measures in online grid operation. An approach to curative grid operation using Online Feedback Optimization is introduced in [27]. OFO is performing supervisory control, recognizing critical system states and activating remedial actions on the same system layer if needed. To further leverage flexibility that is connected to underlying grid layers, i.e. sub-transmission and distribution systems, a hierarchical control structure for an efficient dispatch based on OFO for curative flexibility is introduced in [28, 29]. Different approaches to tuning the proposed control stack are introduced in [30] and [31]. These works illustrate, that the performance of a control stack based on OFO depends on the chosen formulation of the optimization problem and other parameters, e.g. the gain of the controller. The application in safety critical infrastructure like power systems therefore leads to a trade-off between robustness and performance of a proposed control stack. So far previous publications on flexibility quantification, flexibility coordination and feedback-based optimization lack methods to verify a specific control stack with respect to the problem of robustness and performance.
Main Contribution
@@
In order address this gap the main contribution of this work is twofold. First, we introduce an extended hierarchical control architecture for multi-layer flexibility provision in online grid operation. To this end, we incorporate underlying grid layers as flexible resources to support load flow control. Second, we propose a novel method for calculating trajectory sets under uncertainty. This approach verifies the proposed control stack in terms of safety and robustness while quantifying the available flexibility based on the specific implementation and controller tuning. To achieve this, we systematically analyze trajectories resulting from adapting to new load flow set points at the point of common coupling between two grid layers. Based on the theoretical guarantees and previous studies we expect OFO to ensure robust and safe operation of power systems under uncertainty and disturbances by leveraging real-time measurements to iteratively adapt control actions, even in scenarios with limited controllability and incomplete system models.
Closed-Loop Control Architecture
@@
Flexibility Provision
We denote $N$ as the set of buses, $B$ as the set of branches (i.e., lines and transformers), and $F$ as the set of all flexible actors in a grid. Let $\mathbf{v}\in\mathbb{R}^{|N|}$ represent the vector of bus voltages, $\mathbf{s}\in\mathbb{R}^{|B|}$ the vector of active and reactive branch flows, and $\mathbf{p}\in\mathbb{R}^{|F|}$ and $\mathbf{q}\in\mathbb{R}^{|F|}$ the vectors of active and reactive power injections for flexible actors. For a given set point of apparent power flow at the PCC between two grid layers, $\mathbf{s}_{\text{set}}=[p_{\text{set}},q_{\text{set}}]^{T}$, we formulate the following dispatch problem for flexibility:
$\min_{\mathbf{p},\mathbf{q}}$ $\quad\Phi=||p_{\text{set}}-p_{\text{PCC}}||^{2}+||q_{\text{set}}-q_{\text{PCC}}||^{2}$ (1)
@@
This optimization minimizes the Euclidean norm of the difference between the measured apparent power flow, $\mathbf{s}_{\text{PCC}}=[p_{\text{PCC}},q_{\text{PCC}}]^{T}$, and the set point, $\mathbf{s}_{\text{set}}=[p_{\text{set}},q_{\text{set}}]^{T}$, to achieve the desired operating point while satisfying system technical constraints. These constraints, detailed in (1), include limits on bus voltages, $v_{n}\ \forall n\in N$, and branch flows, $s_{i}\ \forall i\in B$. Power limits for available flexibility, such as controllable loads, generation, and other grid layers, are also incorporated as inequality constraints. Since this optimization problem is solved online on the physical system, the AC power flow equations are not explicitly included as equality constraints. Instead, constraint satisfaction is ensured through real-time voltage and branch flow measurements. In the feedback optimization-based approach, the dispatch problem is solved iteratively. At each step, the cost function and constraints are incorporated into a quadratic optimization problem to determine the controller’s next action. The detailed formulation of the proposed OFO controller, including its iterative implementation, is presented in the following section.
Feedback Optimization Based Control
In this section, we introduce the general concept of the OFO controller presented in this work. As stated in the introduction, dispatch problems in the form of (1) are typically solved offline for a model of the system with an estimate of possible disturbances. In contrast, the feedback-based approach iteratively steers the system by directly acquiring the relevant states through online measurements (see Figure 2). In this work we choose projected gradient descent as the underlying optimization algorithm for the controller.
@@
We consider a system $\mathbf{y}\in\mathbb{R}^{p}$ described by its inputs $\mathbf{u}\in\mathbb{R}^{q}$, a function mapping inputs to outputs $h(\mathbf{u})$ and disturbances $\mathbf{d}\in\mathbb{R}^{p}$, where:
$\mathbf{y}=h(\mathbf{u})+\mathbf{d}$ (2)
For our presented use case, this entails a flexibility-providing power system that includes both controllable and non-controllable units, e.g. loads and DERs. The dispatch of flexibility is optimized based on (1) resulting in new operating points for flexible units. After actuation, all units adjust their feed-in or consumption of active and reactive power accordingly.
To ensure stability and convergence to the optimum of the cost function and to allow for an analysis of steady states, we separate the controller’s actions from the fast dynamics of the system [32]. This can be achieved by selecting an appropriate cycle time between the iterations awaiting all relevant faster dynamics of the system such as frequency and voltage transients. In this way we can abstract the differential-algebraic equations of the system describing its physical behavior to their resulting steady states. We express this in the following assumption.
@@
To drive the system to the optimal solution of the underlying optimization problem (1) we approximate the behavior of the system by its steady state map. The proposed controller utilizes an input-output sensitivity map $\nabla h$ that represents the gradient of system output $y$ with respect to control input $u$. The sensitivities are determined by linearly approximating around the smooth power flow manifold for an initial operating point. They include the relevant states, such as bus voltages, branch flows and especially power flow at the PCC. We calculate $\nabla h$ with:
@@
Due to the inclusion of feedback in each step of the controller, $\nabla h$ is computable offline for a fixed initial operating point. We are subsequently able to use this constant approximation in online operation. For the controller to act on the system we define the vector of measurements $y\in\mathbb{R}^{p}$ with
@@
Hierarchical Control Stack
In order to participate in ancillary services, the flexibility from DER at sub-transmission level needs to be aggregated at the TSO-DSO-interface, i.e. the coupling transformer between EHV and HV layer. This can be generalized to the interfaces between flexibility providing and receiving systems, such as further underlying grid layers or Virtual Power Plants (VPPs). Additionally, it is worth noting that the decomposition between the grid layers can also be applied in cases where the transmission and sub-transmission systems are managed by a single system operator to decrease computational complexity in online operation. In the approach presented in this paper, this leads to the following interactions between the supervisory control at TSO-level and the flexibility control at DSO-level. First, the flexibility potential must be estimated by the DSO and made known to the TSO. Second, the TSO needs to request flexibility by sending set points for power flow to the DSO, denoted as $\mathbf{s}_{set}$. To track the requested set points the controller on sub-transmission level needs to update its cost function $\Phi$ upon receiving a new set point (see (1) and (6)). This results in the following sequence of steps for each iteration of the control stack:
@@
    Acquire the current grid state by measurement of power flows and bus voltages.
@@
Robust Trajectory Sets
@@
Feasibility of System States
In this work, we define the term Safety in a control theory sense, i.e. as a system-set property of the power grid, considering its operational constraints, as a dynamical system. A formal definition is provided in [33]. We distinguish this from the term Security, which is more commonly used in context of power systems and refers to the system’s ability to withstand perturbations [34]. To evaluate the safety of system states that result from actuation of a central controller we require a known safe set of such states considering the constraints of the system as a baseline to evaluate against.
@@
In this approach we maximize the load flow for a fixed direction in the solution space, subject to equality constraints $g(\mathbf{x})$ and inequality constraints $h(\mathbf{x})$ of the AC-OPF problem. These include the full grid model with power flow equations in $g(\mathbf{x})$, as well as limits on bus voltages, branch flows, generators, and loads in $h(\mathbf{x})$. The FOR can subsequently be interpreted as a projection of the high-dimensional solution space of (10) onto the two-dimensional PQ-plane at the PCC. This yields the theoretical operating range of the system considering all stationary limits of the detailed grid model. We are therefore able to use the FOR $\mathcal{F}$ to validate individual system states $\mathbf{y}(k)$ that result from the actuation of the controller with respect to these limits. As we choose the inequality constraints $g(\mathbf{x})$ in (1) to be identical to (10) we can make the following assumption for system states resulting from control actions during flexibility provision. We denote the projection of the full system state $\mathbf{y}$ onto the FOR as $\mathbf{y}^{*}$.
@@
Furthermore, we calculate the sensitivities used in the proposed OFO controller based on the same grid model we use to compute $\mathcal{F}$. A full grid model is therefore needed beforehand to perform this evaluation. Here we can distinguish between the operational planning phase and the actual online operation of the controller. In the operational planning phase, we use the full grid model to compute the FOR. During online operation, the controller relies on precomputed sensitivities, reducing the computational complexity. Using Assumption 2, we are able to validate individual system states, system trajectories and consequently trajectory sets that result from the actions of the proposed controller. In the next section we define conditions for the safeness of trajectory sets with respect to the system constraints using the FOR as reference for safe operation.
Feasibility of System Trajectories
@@
For the proposed formulations, safeness implies conditional safeness. Lastly we define unsafe trajectory sets based on constraint satisfaction of the final states of all trajectories $\mathbf{y}(k)\in\mathcal{E}$.
@@
Trajectory Sets under Uncertainty
@@
Using this definition we evaluate a controller stack for a fixed formulation of the optimization problem, a fixed gain $\alpha$ for a set of uncertainties and noise. We describe the transition of system states from $\mathbf{y}_{k-1}$ to $\mathbf{y}_{k}$ by the control actions $\mathbf{u}(k-1)$ and by modeling the individual uncertainties where:
$\mathbf{y}(k)=h^{*}(\mathbf{u})+\omega_{l}+\omega_{y}$ (15)
@@
In cases where we only assume controllability of generation units we consequently consider the current operating points of loads within the system as one of the main disturbances affecting the distribution grid during control of flexibility. We model the variation of the load over time according to [20] for a load covariance matrix $\Sigma_{l}$ with: $\omega_{l}\sim\mathcal{N}(0,\Sigma_{l})$. We differentiate the loads by type (households, industry and commercial) and adjust the entries of the covariance matrix accordingly.
@@
As the proposed control architecture is using direct measurements of bus voltages and branch flows as feedback in the loop it is subject to any noise present in the respective signals. We model the noise term as uniform with $\omega_{y}\sim\mathcal{U}([a,b])$ uncorrelated for all measurements in $y$, in order to specifically evaluate the robustness of the proposed controller with respect to these bounds.
@@
Additionally to the non-modeled dynamics of the system, i.e. uncertainty of load and noisy measurements, we consider a mismatch in model and physical system by introducing bounded errors to the steady-state sensitivity map $\nabla h(u)$. These are described by a bounded uniform distribution $\omega_{s}\sim\mathcal{U}([c,d])$. Similarly to the modeling of the measurement noise, we chose this distribution to provide a bounded approximation of the model mismatch and to be able to evaluate the proposed controller with respect to it. This results in the adjusted sensitivity map $\nabla h^{*}(u)$. Using the uncertainties and their respective distributions, we perform Monte Carlo simulations of the controller converging to reference values on the FOR. This allows for an evaluation of the performance and robustness of the controller in presence of disturbances. We are using the conditions we defined in previous subsections to validate the resulting trajectory sets. Furthermore, we evaluate the iterations needed for convergence to the requested load flows. The results of these case studies are detailed in the next section.
Case Studies
@@
Trajectory Sets at Point of Common Coupling
@@
Influence of Disturbance
This case study examines the behavior of trajectory sets under uncertainty and noisy measurements, focusing on the high-voltage sub-transmission grid described in a previous subsection. The analysis evaluates trajectories that result from controlling load flow at the PCC to the transmission system. Parameter uncertainty, $\omega_{s}$, is modeled as a uniform distribution with bounds of ±$\pm 5\%$, while measurement noise, $\omega_{y}$, is similarly modeled with bounds of ±$\pm 2\%$ on all entries in $y$. First, we consider both load and generation of the system to be fully controllable. The controller is evaluated by performing in total 63.000 simulations of the trajectories with 500 iterations of OFO each under varying disturbances. Based on the resulting system states we can evaluate the density of the different load flows resulting from the control actions at the coupling transformers on $\mathcal{F}$ as shown in Figure 8. We define the density for individual bins on $\mathcal{F}$ based on the number of resulting system states in a bin $n_{ij}$ and the area of the bin $a_{ij}$ with:
$\rho(p,q)=\frac{n_{ij}}{a_{ij}\cdot n_{\text{total}}}$ (16)
@@
The density clearly reflects the general development of the trajectory sets for the controller as shown in Figure 7, indicating that even in when subjected to uncertainties the control actions follow a clear trajectory. It is observable that the density is decreasing for later iterations of the controller. This is explainable by the increasing number of constraints that can become active as the system operates closer to its bounds. Figure 9 illustrates the density of system states lying outside the FOR.
@@
Approximately 40% of all trajectories within the resulting sets contain critical system states. By projecting the overall system states onto the FOR, we identify the limiting constraints under uncertainty. As shown in Figure 9, critical system states are primarily distributed along the lower right section of the FOR, beginning at $\underline{s}_{PCC}=-500\ \text{MW}-j600\ \text{MVAr}$. This distribution is attributed to the controller’s behavior in response to active limiting constraints during operation. Specifically, the constraints imposed by the PCC transformers, as described in the scenario, limit the load flow in this region of the FOR. The observed distribution of critical states suggests that the controller struggles to accommodate the combined effects of parameter uncertainty and transformer constraints in the lower right section of the FOR. A detailed analysis of the sensitivity of the system states to $\omega_{s}$ and $\omega_{y}$ confirms that deviations in measurements significantly affect the load flow trajectory, especially near the constraint-dominated regions. As shown in Figure 10 the final trajectory sets under uncertainty vary significantly from the final set without disturbances.
@@
Especially, the achievable trajectory set of load flows for reactive power provision (top bound of $\mathcal{F}$) to the superimposed system is notably smaller than in the previous case study. This again can be attributed to the density of active constraints in that region of $\mathcal{F}$. As the individual upper bus voltage limits are nearly reached, the noisy measurement leads OFO to perceive these constraints as active. As the noise of the individual sensors is not correlated this leads to a non-convergence of trajectories as OFO is reacting simultaneously to varying perceived constraint violations within the range of the measurement noise $\omega_{y}$. This is further illustrated in Figure 11 showing an exemplary trajectory with and without noisy measurements.
@@
This highlights the difference in influence of measurement uncertainty and sensitivity mismatch as the controller without measurement noise is able to converge even with an augmented sensitivity map. This aligns with previous theoretical considerations guaranteeing convergence of projected gradient descent to an optimal solution for a sufficiently small gain $\alpha$ [22]. To further validate the proposed methods, we analyze a scenario without load controllability and present the corresponding final trajectory set in Figure 12. In this case, the controller attempts to converge to the vertices of the Feasible Operating Region (FOR) with a reduced resolution of $\mathcal{F}$. The reduced availability of controllable power is clearly reflected in the size of the trajectory set, which is significantly smaller compared to the scenario with full controllability of all actuators. The uncertain behavior of the loads introduces additional variability, as depicted in the final trajectory set (gray), diverging from the base scenario without uncertainty. Despite this, the control stack demonstrates robustness as defined in (14), with no observed violations of system limits across all evaluated states. This result highlights the adaptability of the proposed control framework under constrained controllability and uncertain system conditions. While the reduced flexibility imposes stricter limitations on achievable trajectory sets, the controller successfully maintains safe operation by dynamically adjusting to the available flexibility and uncertainties.
@@
Conclusion